# SynSUM – Synthetic Benchmark with Structured and Unstructured Medical Records

### v1

@@ Appendix A Additional example prompts
Figure 7: Example of what would happen if we simply extended the general prompt to the 3629 cases where the patient does not experience any respiratory symptoms. There would be little variation in these notes, as these patients seem to visit the doctor’s office for no reason. It is more realistic to encourage the LLM to generate a note which describes a patient visiting for a non-respiratory complaint, like the prompts shown in Figures 4 and 8.
Appendix B Symptom predictor baselines
@@
Figure 9: Results for the neural-text baseline for the different embedding types. We show F1-scores over the test set for both the normal and compact versions of the notes. Figure best viewed in colour.
B.1 BN-tab
@@
We learn a Bayesian network over the training data, providing the structure over all variables as in Figure 2. For the variables asthma, smoking, hay fever, COPD, season, pneumonia, common cold, fever and self-employed, we learn the conditional probability tables (CPTs) from the training data using maximum likelihood estimation (which comes down to counting co-occurrences of child and parent values for each entry in the CPT). We use the pgmpy library (Ankur Ankan and Abinash Panda, 2015) with a K2 prior as a smoothing strategy to initialize empty CPTs.
@@
As support for learning Noisy-OR distributions is not provided in pgmpy, we learn these parameters with a custom training loop. We formulate the likelihood as in Equation (2), and learn the parameters $p_{i}$ in Equations (3) through (6) for the variables dyspnea, cough, pain and nasal through maximum likelihood estimation by iterating over the train set for 10 epochs, using an Adam optimizer with a batch size of 50, a learning rate of 0.01 and random initialization of each parameter. To integrate the learned Noisy-OR distributions in the Bayesian network, we turn them into fully specified CPTs. To obtain these, we simply evaluate Equation (2) for all possible combinations of child and parent values. While this results in large and inefficient CPTs, the automated inference engine bulit into pgmpy library does not support Noisy-OR distributions directly. Note that both versions of the conditional distribution are equivalent, so we do not incur a loss in precision.
@@
Similarly, the coefficients in the logistic regression model for antibiotics and the Poisson regression model for #days are learned using maximum likelihood estimation over the training set. The likelihood is expressed as in Equation (7) and Equations (8) and (9) respectively, with learnable parameters in place of each coefficient. We iterate over the train set for 15 epochs, again using an Adam optimizer with a batch size of 50, a learning rate of 0.01 and random initialization of each parameter. Finally, we turn the logistic regression and Poisson regression models into CPTs by evaluating Equations (7), (8) and (9) for all combinations of parent and child values. For the variable #days, we needed to turn each discrete number of days into a category, because pgmpy only provides automated inference for Bayesian networks consisting of exclusively categorical variables. This results in a large CPT containing one row per possible number of days, which range from 0 to 15 in our training dataset, and one column for each combination of the 7 parent variables. To allow for a possible larger maximum number of days in the test set, we create a category $\geq\text{{15 days}}$, which is defined as one minus the summed probability of all other days.
@@ B.3 Neural-text
The pretrained BioLORD encoder (Remy, Demuynck, and Demeester, 2024) was obtained through the huggingface library. The encoder outputs 768-dimensional sentence embeddings. Since the full text did not fit into the context window, we embedded each sentence separately, and then combined them using our strategies outlined in the main text. To split the text into sentences, we used the nltk package. The settings hist, phys and mean all result in a text embedding of 768 dimensions, while the setting concat results in a text embedding of 2*768 dimensions.
@@
These embeddings are fed into a linear layer with 256 neurons, which is then transformed into a single output neuron, followed by a Sigmoid activation. For the classifiers that predict fever, three output neurons followed by a Softmax activation are used instead, one for each class. While the embeddings remain fixed, we learn the parameters in the hidden and output layers using cross-entropy as a loss function over the training set. We train a separate classifier for each symptom, setting and difficulty of the text (normal vs. compact). For the binary symptoms, we train for $15$ epochs using the Adam optimizer with a batch size of $100$, a learning rate of $0.001$ and weight_decay set to $51\mathrm{e}-$ 5. The classifier for $fever$ tended to collapse more easily, which is why we train it for $30$ epochs with a lower learning rate of $45\mathrm{e}-$ 4 instead. These hyperparameters were obtained using a mix of manual tuning and grid search with 5-fold cross validation over the training set.
Results for the different embedding types, which are reported numerically in Table 7, are shown in Figure 9.
B.4 Neural-text-tab
@@
For each evidence setting, we select the relevant set of tabular features and transform them into a numerical representation. We use a one-hot encoding for the categorical (binary or multiclass) features, and normalize the #days feature using the StandardScaler from sklearn. This tabular feature representation is then concatenated with the text representation we obtained in the previous baseline. Both are fed into the same architecture described in Section B.4, adapting the dimension of the input layer accordingly. For example, for the dyspnea classifier in the evidence setting all, the input dimension becomes $768+17$. We use the same hyperparameters as in the neural-text baseline to ensure a fair comparison. All other training details remain the same.

### v2

@@ Appendix D Additional example prompts
Appendix E Expert evaluation
We picked a random subset of $30$ generated notes and show them to $5$ general practitioners, in a random order. All evaluators got to see the same $30$ notes, together with the prompts that were used to generate them. All evaluators received detailed instructions on what was expected of them in the form of a PDF [^1] https://github.com/prabaey/SynSUM/blob/main/eval/Instructions˙clinical˙evaluation.pdf, which was orally explained by the researchers. The researchers then sat together with each evaluator separately to complete three example notes (different from the $30$ notes that were to be evaluated). Afterwards, the evaluators were asked to rate the notes in their own time, without the researchers’ involvement. Since it is infeasible to evaluate the whole dataset of $10{,}000$ notes, we opted for a small subset of $30$ notes, each going through a relatively extensive evaluation process that considered various measures of quality (evaluators took around $5-10$ minutes to evaluate each note). We decided to show all $5$ evaluators the same set of $30$ notes, to get a broader range of expert opinions in the evaluation of each note. This also allowed for the calculation of inter-annotator agreement.
We now provide further details on each dimension along which we evaluated the notes and the specific meaning assigned to each rating. The aspects of consistency, realism and clinical accuracy are only evaluated based on the normal note, while the quality of the compact note is evaluated using the last two dimensions (content and readability).
E.1 Evaluation aspects
Consistency
We subdivided the prompt into four different sections, and asked the evaluators to assign penalties for each section. A penalty was assigned if the requested information in that section was incorrectly mentioned in the note (e.g. a particular symptom was said to be present in the patient, when the prompt particularly requested the symptom to be absent), or if the requested information was absent from the note (e.g. a symptom descriptor is not mentioned in the text). The four parts of the prompt were as follows: (i) the symptoms to mention (can be present or absent), (ii) the symptoms not to mention, (iii) the underlying health conditions, and (iv) the additional instructions.
As explained in Section 2.2, around one third of the notes were generated using a second type of prompt, where the LLM is told that the patient does not suffer from any respiratory symptoms or underlying health conditions. Following the prompt in Figure 6, there are three parts of the prompt which can be violated (leading to penalties): (i) the respiratory symptoms, which the patient does not have, (ii) the underlying conditions, which the patient does not have, and (iii) the additional instructions. In our random set of $30$ notes, $20$ notes belonged to the first type, and $10$ to the second type.
Once penalties were assigned, we summed them into a total number of penalties, and converted these into scores from $1$ to $5$. Notes with no penalties get a perfect score of $5$, one penalty corresponds to $4$, two penalties to $3$, three penalties to $2$ and more than three penalties to $1$.
Realism
The LLM is allowed to invent context and details in light of the information it receives in the prompt, but this must be realistic and relevant to the symptoms experienced by the patient. While some clinical facts might not seem technically incorrect, one might not expect to see them in the note, or it might be unlikely that they would be written down by a real physician. For example, if a patient has a runny nose and no other complaints, most clinicians would not check for abnormalities in lung capacity. Another example is asking whether the patient has recently traveled to an exotic destination because they have a cough.
We ask to score realism of the “history” section using the ratings below. The evaluators are specifically instructed to take into account the information mentioned in the prompt.
5 – All pieces of additional context and details (i.e. outside of the symptoms and background provided in the prompt) are realistic and seem like they belong in the note.
4 – There are one or two pieces of additional context or details that I would not have mentioned as a physician, or that do not seem relevant (even though they do seem like they belong).
3 – There are one or two pieces of additional context or details that do not seem like they belong in the note, or do not seem relevant, given the symptoms and background provided in the prompt.
2 – There are multiple pieces of additional context or details that do not seem like they belong in the note, or do not seem relevant, given the symptoms and background provided in the prompt.
1 – (Almost) all of the additional context is nonsensical given the symptoms and background provided in the prompt.
We ask to score realism of the “physical examination” section using the ratings below. The evaluators are specifically instructed to take into account the information mentioned in the “history” section of the note.
5 – All elements in the physical examination are things I would check, given the history and symptoms of the patient, and no important elements are missing.
4 – There are one or two elements in the physical examination that I probably would not have checked, given the history and symptoms of the patient, but I could see it happen. Some minor elements might be missing, but nothing major.
3 – There are one or two elements in the physical examination that I would not have checked, or some important elements are missing, given the history and symptoms of the patient.
2 – There are multiple elements in the physical examination that make no sense given the history and symptoms of the patient, or many important elements are missing.
1 – The physical examination portion of the note seems totally unrealistic.
Clinical accuracy
While the previous section talks about evaluating the realism of the presence of all examinations described in the “physical examination” section, here we talk about evaluating the clinical accuracy of these findings. Clinical inaccuracies may depend on the context, like physical findings which are not congruent with the history and symptoms of the patient. For example, if the “history” portion mentions that the patient has a no fever, then this should not be contradicted in the “physical examination” portion with a temperature of 39°C. Clinical inaccuracies may also stand alone. For example, a blood pressure reading of 20/10 mm Hg is impossible to encounter in any patient.
We ask the evaluators to score clinical accuracy of the findings that are mentioned in the “physical examination” portion of the note, using the ratings below.
5 – There are no mistakes, all reported clinical information is plausible in light of the patient’s symptoms and history.
4 – There are one or two minor mistakes, or some details seem less plausible in light of the patient’s symptoms and history, while the overall picture painted by the note is still correct.
3 – There are more than two minor mistakes, or multiple details which seem implausible in light of the patient’s symptoms and history, but no major inaccuracies.
2 – There is a major mistake (on top of possibly some minor ones), or many details seem implausible given the patient’s symptoms and history.
1 – There are multiple major mistakes and many details seem totally implausible given the patient’s symptoms and history.
Quality of compact version
While all the previous evaluations concerned the original note, here we evaluate the quality of the compact version of the note.
The content of the compact version should convey the same information as the original text, albeit in a shorter format. This is evaluated jointly for “history” and “physical examination” using the scoring system below.
5 – The compact version conveys the exact same information as the original text.
4 – The compact version conveys all key points of the original text, leaving out some details here and there.
3 – The compact version conveys some of the key points of the original text, but misses some as well.
2 – The compact version conveys some of the same information as the original text, but misses many key points.
1 – The compact version does not convey the same information as the original text, leaving out almost all key points.
While we purposefully want these notes to be harder to read and understand for both humans and machines, mimicking the complexity of some real clinical notes, the use of abbreviations should not be excessive. We evaluate readability jointly for “history” and “physical examination”, using the scoring system below.
5 – The compact version seems understandable without seeing the original.
4 – The compact version seems mostly understandable without seeing the original, though there are some abbreviations that I would not immediately understand.
3 – Some parts of the compact version seem understandable without seeing the original, but other parts are not. There are some abbreviations that seem far-fetched or are used incorrectly (i.e. these are known to refer to other clinical terms than the way they are used in the text).
2 – Many parts of the compact version would not be understandable without seeing the original. Many abbreviations seem far-fetched or are used incorrectly (i.e. these are known to refer to other clinical terms than the way they are used in the text).
1 – The compact version is impossible to understand without seeing the original.
E.2 Results
Extended results for consistency
We list the total number of penalties assigned by all evaluators to each of the prompt sections. For the 20 notes belonging to the first prompt type, 41 penalties were assigned in total for section (iv) of the prompt, which are the additional instructions. Sections (i), (ii), and (iii), which describe the symptoms and underlying health conditions of the patients, received no penalties at all. For the 10 notes belonging to the second prompt type, 1 penalty was assigned for part (i), 2 penalties for part (ii) and another 2 penalties for part (iii).
From this, we can conclude that a large majority of the inconsistencies arise from a violation of the additional instructions (usually by inventing additional symptoms), while the key symptom information included in the prompt was still conveyed correctly in the note. High consistency between the tabular variables and the concepts mentioned in the text makes our dataset a reliable resource for information extraction tasks.
Inter-Annotator Agreement
We calculate Krippendorff’s alpha with ordinal distance measure to get an idea of the inter-annotator agreement in our study. We get the following results:
  • Consistency: 0.44
  • Realism (history): 0.25
  • Realism (physical examination): 0.32
  • Clinical accuracy: 0.21
  • Content of compact version: -0.02
  • Readability of compact version: 0.36
With an alpha of 1 indicating perfect agreement, we can say that there is some agreement between the 5 evaluators. Especially for consistency, readability and realism of the physical examination, the agreement is fair. Only the ratings for content of the compact version shows no indication of agreement. We hypothesize that this is at least partially due to the high scores assigned by all evaluators (mostly 4 and 5), with a deviation from 5 being assigned to random chance rather than to a common opinion of the evaluators. The lower standard deviations in Table 1 show that the evaluators consistently assign high scores to all aspects of the evaluation, even if they don’t agree on which particular notes deviate from these higher scores.
Appendix F Symptom predictor baselines
@@
Figure 9: Results for the neural-text baseline for the different embedding types. We show F1-scores over the test set for both the normal and compact versions of the notes. Figure best viewed in color.
F.1 BN-tab
@@
We learn a Bayesian network over the training data, providing the structure over all variables as in Figure 1. For the variables asthma, smoking, hay fever, COPD, season, pneumonia, common cold, fever and self-employed, we learn the conditional probability tables (CPTs) from the training data using maximum likelihood estimation (which comes down to counting co-occurrences of child and parent values for each entry in the CPT). We use the pgmpy library (Ankur Ankan and Abinash Panda 2015) with a K2 prior as a smoothing strategy to initialize empty CPTs.
@@
As support for learning Noisy-OR distributions is not provided in pgmpy, we learn these parameters with a custom training loop. We formulate the likelihood as in Equation (2), and learn the parameters $p_{i}$ in Equations (3) through (6) for the variables dyspnea, cough, pain and nasal through maximum likelihood estimation by iterating over the train set for 10 epochs, using an Adam optimizer with a batch size of 50, a learning rate of 0.01 and random initialization of each parameter. To integrate the learned Noisy-OR distributions in the Bayesian network, we turn them into fully specified CPTs. To obtain these, we simply evaluate Equation (2) for all possible combinations of child and parent values. While this results in large and inefficient CPTs, the automated inference engine built into pgmpy library does not support Noisy-OR distributions directly. Note that both versions of the conditional distribution are equivalent, so we do not incur a loss in precision.
@@
Similarly, the coefficients in the logistic regression model for antibiotics and the Poisson regression model for #days are learned using maximum likelihood estimation over the training set of $8{,}000$ examples. The likelihood is expressed as in Equation (7) and Equations (8) and (9) respectively, with learnable parameters in place of each coefficient. We iterate over the train set for 15 epochs, again using an Adam optimizer with a batch size of 50, a learning rate of 0.01 and random initialization of each parameter. Finally, we turn the logistic regression and Poisson regression models into CPTs by evaluating Equations (7), (8) and (9) for all combinations of parent and child values. For the variable #days, we needed to turn each discrete number of days into a category, because pgmpy only provides automated inference for Bayesian networks consisting of exclusively categorical variables. This results in a large CPT containing one row per possible number of days, which range from 0 to 15 in our training dataset, and one column for each combination of the 7 parent variables. To allow for a possible larger maximum number of days in the test set, we create a category $\geq\text{{15 days}}$, which is defined as one minus the summed probability of all other days.
@@ F.3 Neural-text
The pretrained BioLORD encoder (Remy, Demuynck, and Demeester 2024) was obtained through the huggingface library. The encoder outputs 768-dimensional sentence embeddings. Since the full text did not fit into the context window, we embedded each sentence separately, and then combined them using our strategies outlined in the main text. To split the text into sentences, we used the nltk package. The settings hist, phys and mean all result in a text embedding of 768 dimensions, while the setting concat results in a text embedding of 2*768 dimensions.
@@
These embeddings are fed into a linear layer with 256 neurons, followed by a ReLU activation. The hidden state is then transformed into a single output neuron, followed by a Sigmoid activation. For the classifiers that predict fever, three output neurons followed by a Softmax activation are used instead, one for each class. While the embeddings remain fixed, we learn the parameters in the hidden and output layers using cross-entropy as a loss function over the training set. We train a separate classifier for each symptom, setting and difficulty of the text (normal vs. compact). For the binary symptoms, we train for $15$ epochs using the Adam optimizer with a batch size of $100$, a learning rate of $0.001$ and weight_decay set to $51\mathrm{e}-$ 5. The classifier for $fever$ tended to collapse more easily, which is why we train it for $30$ epochs with a lower learning rate of $45\mathrm{e}-$ 4 instead. These hyperparameters were obtained using a mix of manual tuning and grid search with 5-fold cross validation over the training set.
Results for the different embedding types are reported numerically in Table 8, and shown visually in Figure 9. There is a significant gap in performance between the hist and phys settings for the symptoms cough, pain and fever. This makes sense, as the “history” section of the note outlines the symptoms experienced by the patient more clearly. The performance difference between the mean and concat settings is usually small.While the score for hist comes close to those for mean and concat, the latter usually still outperform the former for the normal notes, showing that there is some complementary information in the “history” and “physical examination” portions. For the compact notes, including the “physical examination” portion seems to confuse the model for the symptoms cough and pain, since the hist setting outperforms mean and concat there.
dyspnea cough pain nasal fever
normal
- hist 0.9399 0.9699 0.8310 0.9538 0.9028
- phys 0.9035 0.7948 0.6871 0.9526 0.8580
- mean 0.9660 0.9595 0.8415 0.9602 0.9125
- concat 0.9635 0.9660 0.8313 0.9606 0.9001
compact
- hist 0.9353 0.9557 0.8008 0.9522 0.8978
- phys 0.8747 0.7655 0.6492 0.9544 0.8324
- mean 0.9383 0.9480 0.7828 0.9583 0.9073
- concat 0.9358 0.9443 0.7849 0.9598 0.8964
Table 8: Results for the neural-text baseline using different embedding types at the input. We report F1-score over the test set for both the normal and compact versions of the notes. The best results per symptoms and per note version is shown in bold.
F.4 Neural-text-tab
@@
For each evidence setting, we select the relevant set of tabular features and transform them into a numerical representation. We use a one-hot encoding for the categorical (binary or multiclass) features, and normalize the #days feature using the StandardScaler from sklearn. This tabular feature representation is then concatenated with the text representation we obtained in the previous baseline. Both are fed into the same architecture described in Section F.4, adapting the dimension of the input layer accordingly. For example, for the dyspnea classifier in the evidence setting all, the input dimension becomes $768+17$. We use the same hyperparameters as in the neural-text baseline to ensure a fair comparison. All other training details remain the same.